# Chika Maduabuchi, Ericmoore Jossou, Matteo Bucci

### v3

@@ Abstract
Purpose: High-speed video (HSV) phase detection (PD) segmentation is vital in nuclear reactors, chemical processing, and electronics cooling for detecting vapor, liquid, and microlayer phases. Traditional segmentation models face pixel-level accuracy and generalization issues in multimodal data. MSEG-VCUQ introduces VideoSAM, a hybrid framework leveraging convolutional neural networks (CNNs) and transformer-based vision models to enhance segmentation accuracy and generalizability across complex multimodal PD tasks.
Methods: VideoSAM combines U-Net CNN and the Segment Anything Model (SAM) for advanced feature extraction and segmentation across diverse HSV PD modalities, spanning fluids like water, FC-72, nitrogen, and argon under varied heat flux conditions. The framework also incorporates uncertainty quantification (UQ) to assess pixel-based discretization errors, delivering reliable metrics such as contact line density and dry area fraction under versatile modalities.
Results: VideoSAM outperforms SAM and modality-specific CNN models in segmentation accuracy, excelling in environments with complex phase boundaries, overlapping bubbles, and dynamic liquid-vapor interactions. Its hybrid architecture supports cross-dataset generalization, adapting effectively to varying modalities. The UQ module provides accurate error estimates, enhancing the reliability of segmentation outputs for advanced HSV PD research.
Conclusion: MSEG-VCUQ, via VideoSAM, offers a robust solution for HSV PD segmentation, addressing previous limitations with advanced deep learning and UQ techniques. The open-source datasets and tools introduced enable scalable, precise, and adaptable segmentation for multimodal PD datasets, supporting advancements in HSV analysis and autonomous experimentation. The codes and data used for this paper are publicly available at: https://github.com/chikap421/mseg_vcuq.
Keywords Phase Detection Segmentation  $\cdot$ High-Speed Video Analysis  $\cdot$ Convolutional Neural Networks  $\cdot$ Vision Foundation Models  $\cdot$ Uncertainty Quantification
@@ 1 Introduction
Boiling heat transfer is a complex phenomenon integral to various industrial and engineering applications, including nuclear reactors, electronics cooling, and chemical processing systems, where efficient thermal management is paramount [1, 2]. This process is governed by intricate mechanisms such as evaporation, quenching, and single-phase convection, all contributing to the overall heat flux [3]. Predictive mechanistic models aim to quantify these contributions using key parameters like nucleation site density, bubble departure diameter, contact line density, and dry area fraction. These metrics are crucial for understanding boiling dynamics but depend heavily on high-resolution data obtained through advanced diagnostics such as infrared thermography [2, 4] and high-speed video (HSV) phase detection (PD) imaging [5, 6].
@@
PD images, such as those shown in Figure 1, provide crucial insights by differentiating between liquid, vapor, and microlayer phases on a boiling surface [5]. Yet, traditional methods for analyzing these images, whether manual or semi-automated, require significant time and expertise, limiting scalability for larger multimodal datasets. Convolutional neural networks (CNNs), especially U-Net [7, 8], have successfully segmented two-phase flow (TPF) images by learning complex features in data, thus providing a promising solution to automate HSV PD segmentation. Unlike the shadowgraphy images typically used for TPF studies (Figure 1), PD images feature complex bubble structures and diverse liquid-vapor interactions. These unique attributes create challenges for accurate segmentation, particularly in identifying the bubble footprints and contact lines.
@@
Figure 1: (a) Sample of Front-Lit Shadowgraphy images from two-phase flow. (b) Sample Phase-Detection Images used in this study.
@@
HSV PD data offers a valuable lens into the boiling process by delineating liquid, vapor, and microlayer phases. However, these HSV datasets’ manual or semi-automated analysis can be labor-intensive and subject to subjectivity, limiting scalability for large multimodal datasets. Traditional segmentation techniques, including edge detection and thresholding, struggle with overlapping bubbles and dynamic flow patterns, often requiring manual adjustments across experiments. Recently, CNN architectures like U-Net have become the standard for HSV tasks due to their ability to learn intricate visual patterns and deliver precise segmentation results [9, 10, 11]. Nevertheless, these CNN architectures, though powerful, exhibit limited generalizability when applied to multimodal datasets comprising varying experimental conditions, posing a significant barrier to their adoption in scientific PD tasks [12, 13, 14].
@@
This limitation is pronounced in HSV PD analysis, where dynamic physical processes and domain-specific visual characteristics demand flexible yet robust segmentation models. The emergence of vision foundation models, like the Segment Anything Model (SAM), has demonstrated promising generalization capabilities across various segmentation tasks by leveraging transformer-based architectures [15]. Despite their potential, adapting these models to scientific HSV PD tasks remains underexplored, with most applications focused on natural images rather than PD in boiling experiments [16].
@@ 2 Related Works
Developing image segmentation techniques for two-phase flow (TPF) analysis showcases a relentless pursuit of precision, evolving from classical image processing tools to sophisticated deep learning models and, more recently, to scalable vision foundation architectures. Though effective under controlled conditions, traditional algorithms faltered when confronted with complex flow patterns, overlapping interfaces, and fluctuating thermal dynamics. These early methods, including edge detection and threshold-based segmentation, lacked generalizability across varying experimental setups, requiring manual adjustments for each scenario. This inconsistency underscored the need for more flexible and adaptive frameworks.
Introducing deep learning architectures, such as CNNs like U-Net and Mask R-CNN, significantly addressed these challenges. These models captured intricate phase boundaries and dynamic behaviors in TPF images. However, even CNN-based approaches encountered limitations in PD scenarios, demanding pixel-level precision to effectively distinguish vapor, liquid, and microlayer phases. Transfer learning emerged as a promising solution, enabling models trained on analogous domains—such as biological imaging—to enhance performance on TPF datasets. Yet, the task-specific nature of these CNNs restricted their adaptability, especially under new experimental conditions or with previously unseen modalities, leaving room for further exploration in multimodal HSV PD datasets.
In response to these gaps, large vision foundation models like the Segment Anything Model (SAM) have gained traction due to their potential to generalize across diverse datasets with minimal fine-tuning. These models, built on extensive pretraining across various domains, offer a new horizon in segmentation, particularly in complex scientific applications like PD. Our discussion in this section synthesizes insights from prior work [17, 18], providing a nuanced review of the landscape. We examine the evolution from classical image processing to CNNs, exploring their strengths and weaknesses, and then pivot to the emerging role of large vision models. This integrated review highlights the limitations of existing methods and underscores the importance of scalable, adaptable frameworks like SAM. By doing so, we set the stage for our contribution: leveraging transfer learning, robust uncertainty quantification, and open-source datasets to bridge the gaps in PD segmentation and advance HSV PD segmentation research.
2.1 Traditional Image Processing Methods in Two-Phase Flow Analysis
Before adopting deep learning, various image-processing algorithms were employed to track bubbles in TPF scenarios. Early approaches focused on classical computer vision techniques, including Canny edge detection enhanced with Gaussian smoothing to mitigate noise. Wenyin et al. [19] leveraged these methods to improve the segmentation of gas-liquid interfaces from HSV data. Their approach utilized gradient non-maximum suppression and dual thresholding, extracting bubble contours with notable precision. However, these algorithms encountered challenges when applied to complex scenarios involving overlapping bubble boundaries, highlighting the need for further improvements.
Building on these earlier efforts, Paz et al. [20] proposed machine vision algorithms that extended segmentation capabilities for bubbles in subcooled TPF boiling phenomena. Their work incorporated side and front views of the boiling process, improving the adaptability of segmentation across different experimental setups. Dynamic thresholding and edge detection allowed the extraction of bubble outlines and trajectories, facilitating more detailed flow analysis. However, setup calibration and parameterization remained key challenges, requiring manual tuning to achieve optimal performance under diverse conditions.
Zhou and Niu [21] advanced bubble detection with a multi-frame image processing algorithm in another significant contribution. Their approach utilized binarization and a predictor-corrector method to handle overlapping bubble images better. This method significantly enhanced bubble size and velocity measurements in experiments featuring dense bubbly plumes. Despite these improvements, the algorithm struggled to distinguish individual bubbles when significant overlap occurred, treating merged bubbles as single entities. This limitation underscored the need for more adaptive segmentation methods capable of performing well in varied flow conditions.
Further refinement came from Richenderfer et al. [1], who developed an experimental methodology using HSV and infrared thermometry to measure key boiling parameters, including nucleation site density and bubble dynamics. Their method incorporated a combination of gray threshold filters and watershed segmentation to partition heat flux accurately, providing insights into the contribution of microlayer evaporation. Although effective, traditional thresholding techniques faced limitations under high heat flux conditions, where they often failed to segment bubbles accurately, prompting the need for more advanced methodologies.
These early efforts [22, 23, 24] illustrate the potential of traditional image processing methods for segmenting bubble behavior in HSV data. They also expose their inherent shortcomings, particularly in the face of variable lighting conditions, overlapping bubbles, and fluid-specific properties. The absence of a universally applicable segmentation framework necessitated the shift towards deep learning architectures, capable of adapting to the complex, dynamic nature of boiling phenomena through learned features and robust generalization.
2.2 Deep Learning Architectures for HSV Two-Phase Flow Segmentation
Traditional CNN-based segmentation tools such as U-Net [25] and Mask R-CNN [26] have become the backbone of bubble segmentation. These models were initially employed to detect objects such as bubbles, droplets, and particles within heated or cooled fluids, often under varying conditions. Studies by Passoni et al. [14] and Malakhov et al. [8] highlight the success of these CNNs, with U-Net excelling in detecting smaller structures but struggling with generalization when faced with unseen flow conditions. While computationally heavier, Mask R-CNN achieved commendable instance segmentation by leveraging the COCO dataset pre-trained weights [27], allowing more detailed shape extraction of overlapping bubbles.
Transfer learning in these models has emerged as a critical step to overcome the limitations of small datasets. For instance, Kim and Park [28] enhanced Mask R-CNN’s capabilities by pre-training on biological cell imagery, demonstrating that domain alignment plays a pivotal role in model performance. They achieved an average precision of 0.981, reducing mask extraction times compared to manual techniques. Similarly, Soibam et al. [7] utilized transfer learning to develop a robust YOLOv7 model tailored for subcooled boiling scenarios. Their work demonstrates that combining deep learning with transfer learning enables models to achieve higher Intersection over Union (IoU) scores (88%) than traditional algorithms.
Despite these advancements, several models still need help. Seong et al. [12] identified that traditional U-Net struggled to capture intricate phase boundaries while effectively segmenting specific bubbles, limiting its utility across diverse conditions. Ravichandran et al. [13] addressed this by developing a 3D U-Net that incorporated infrared (IR) thermometry data for more accurate detection of dry patches. However, their model required extensive customization for new experimental setups, reducing scalability.
Suh et al. [29] introduced the Vision-Inspired Online Nuclei Tracking System (VISION-iT), which integrated Mask R-CNN with post-processing tools to capture nucleation events in HSV data. Although this system offers precise tracking, it demands significant fine-tuning to adapt to new datasets, hindering its general applicability across diverse experimental frameworks.
2.3 Vision Foundation Models for Segmentation
Recent strides in computer vision have led to the development of large vision foundation models, such as SAM, designed to enhance generalization and performance across diverse segmentation tasks. These models surpass traditional CNN-based approaches by leveraging extensive pretraining on large-scale datasets, capturing global and local features, and excelling in learning long-range dependencies critical for complex segmentation tasks [30]. Unlike CNNs, which primarily focus on local spatial features, models like Swin Transformer introduce hierarchical attention mechanisms that enhance feature extraction across multiple scales [31]. Prompt-based learning frameworks like SAM offer rapid adaptation to new segmentation challenges, ensuring flexibility and robustness across varied domains with minimal fine-tuning [32]. This architectural adaptability makes them superior for diverse segmentation tasks [33].
SAM, for example, demonstrates versatility by excelling in natural image segmentation and achieving breakthroughs in medical imaging [34, 35, 36], remote sensing [37, 38, 39], and video tracking [40, 41, 42]. Other prominent models include SEEM for multimodal segmentation [43], Mask2Former with masked attention mechanisms for universal segmentation [33], HRNet with high-resolution representations [30], and Swin Transformer for hierarchical feature extraction [31]. These innovations address scalability limitations in smaller CNN models, paving the way for their integration into more complex scientific tasks like phase detection and fluid dynamics.
2.4 Existing Datasets in Boiling Phenomena
Traditional datasets, such as the Boiling Dataset, [44, 45] focus on broader TPF phenomena, targeting tasks like classification and flow regime identification. However, these datasets fall short regarding PD tasks that demand pixel-level precision. TPF datasets often lack high-resolution frame-mask pairs, limiting their applicability in training advanced deep-learning models for precise segmentation, such as those used in HSV imagery. This gap hinders accurately segmenting vapor, liquid, and microlayer phases.
To address these challenges, we introduce a novel PD dataset tailored for HSV segmentation. This dataset bridges the gap by offering high-resolution frame-mask pairs with detailed annotations for liquid, vapor, and microlayer phases. The open-source nature of this dataset encourages further research, enabling the fine-tuning of state-of-the-art models for new experimental setups and various fluid types. Our contribution advances PD segmentation and provides a benchmark that fosters innovation in both scientific and industrial applications of HSV PD segmentation.
2.5 Uncertainty Quantification in Boiling Phenomena Segmentation
Several studies emphasize the necessity of robust UQ techniques in fluid dynamics. Ravichandran et al. [13] explored UQ in the context of deep learning models for boiling heat transfer, highlighting how uncertainties in segmented data can skew estimations of thermal performance, ultimately influencing system-level insights. Seong et al. [12] further reinforced the importance of UQ, particularly in HSV datasets, where overlapping objects like bubbles pose significant challenges to segmentation accuracy. Without proper uncertainty estimation, these models are prone to delivering erroneous conclusions, especially in PD tasks that require pixel-level precision. These findings align with earlier works that called for greater emphasis on UQ to enhance robustness in medical image segmentation [46, 47], radiotherapy [48], and bubble segmentation applications [49].
2.6 Research Gaps in HSV PD Segmentation
Despite notable advancements in CNN-based models for TPF segmentation, a fundamental gap persists in their applicability to multimodal HSV PD segmentation. PD requires finer distinctions between vapor, liquid, and microlayer phases, which demand pixel-level precision. Traditional TPF segmentation frameworks, including popular CNN architectures like U-Net and Mask R-CNN, often focus on object-level boundaries, limiting their ability to capture the subtle transitions and overlapping phases characteristic of PD. This constraint becomes particularly pronounced in dynamic multimodal scenarios involving fluctuating heat flux conditions and complex experimental setups, where precise delineation of phases directly impacts the accuracy of boiling heat transfer models. Our research bridges this gap by introducing PD-specific segmentation that leverages transfer learning from biological cell domains, ensuring custom models are better adapted to PD images. Transfer learning allows efficient training on limited datasets by utilizing knowledge from analogous domains, such as cell imaging, where precise segmentation is similarly required. This adaptation enhances the custom U-Net model’s ability to handle the complex geometry and variability of boiling dynamics under different modalities, thus overcoming limitations seen in earlier work that focused on simpler TPF cases.
A significant innovation in our approach is the integration of UQ within the segmentation pipeline. UQ addresses the inherent errors in pixel-based measurements, providing robust metrics for parameters like contact line density and dry area fraction. Our framework systematically quantifies discretization errors through advanced open-source scripts and ensures reliable outputs across diverse modalities. This structured UQ approach offers a much-needed refinement in error estimation, addressing gaps identified in prior studies where uncertainty in predictions led to unreliable insights into phase transitions and heat transfer mechanisms.
Moreover, we address a critical bottleneck in the field: the limited availability of open-source datasets tailored for PD tasks. Existing datasets predominantly target broader TPF segmentation or classification, leaving a gap in resources necessary for advancing PD-specific research. Our work introduces open-source datasets to facilitate cross-dataset adaptability and model development, empowering researchers to test and fine-tune their models under multimodal conditions. Additionally, by extending the segmentation task to large vision foundation models, such as SAM, our research bridges the scalability gap, ensuring these models generalize effectively across different modalities and experimental conditions. Fine-tuning SAM on our novel PD datasets provides a flexible and scalable solution, enabling seamless adaptation to new scenarios where custom U-Net CNN models were limited by specificity.
Our comprehensive framework—encompassing segmentation, uncertainty quantification, dataset development, and cross-dataset adaptability—sets the foundation for further exploration and development in this domain. By providing open-source tools and resources, we advance the scientific understanding of PD segmentation and foster future innovation, establishing a versatile platform for addressing the intricacies of boiling heat transfer across diverse research contexts.

### v4

# MSEG-VCUQ: Multimodal SEGmentation with Enhanced Vision Foundation Models, Convolutional Neural Networks, and Uncertainty Quantification for High-Speed Video Phase Detection Data

@@ Abstract
High-speed video (HSV) phase detection (PD) segmentation is crucial for monitoring vapor, liquid, and microlayer phases in industrial processes. While CNN-based models like U-Net have shown success in simplified shadowgraphy-based two-phase flow (TPF) analysis, their application to complex HSV PD tasks remains unexplored, and vision foundation models (VFMs) have yet to address the complexities of either shadowgraphy-based or PD TPF video segmentation. Existing uncertainty quantification (UQ) methods lack pixel-level reliability for critical metrics like contact line density and dry area fraction, and the absence of large-scale, multimodal experimental datasets tailored to PD segmentation further impedes progress. To address these gaps, we propose MSEG-VCUQ. This hybrid framework integrates U-Net CNNs with the transformer-based Segment Anything Model (SAM) to achieve enhanced segmentation accuracy and cross-modality generalization. Our approach incorporates systematic UQ for robust error assessment and introduces the first open-source multimodal HSV PD datasets. Empirical results demonstrate that MSEG-VCUQ outperforms baseline CNNs and VFMs, enabling scalable and reliable PD segmentation for real-world boiling dynamics.
Machine Learning, ICML
@@ 1 Introduction
Boiling heat transfer is a complex phenomenon integral to various industrial and engineering applications, including nuclear reactors, electronics cooling, and chemical processing systems, where efficient thermal management is paramount (Richenderfer et al., 2018; Chi Wang et al., 2024). This process is governed by intricate mechanisms such as evaporation, quenching, and single-phase convection, all contributing to the overall heat flux (Bucci et al., 2016). Predictive mechanistic models aim to quantify these contributions using key parameters like nucleation site density, bubble departure diameter, contact line density, and dry area fraction. These metrics are crucial for understanding boiling dynamics but depend heavily on high-resolution data obtained through advanced diagnostics such as infrared thermography (Chi Wang et al., 2024; Wang et al., 2023) and high-speed video (HSV) phase detection (PD) imaging (Kossolapov et al., 2021, 2024).
@@
PD images, such as those shown in Figure 1b, provide crucial insights by differentiating between liquid, vapor, and microlayer phases on a boiling surface (Kossolapov et al., 2021). Yet, traditional methods for analyzing these images, whether manual or semi-automated, require significant time and expertise, limiting scalability for larger multimodal datasets. Convolutional neural networks (CNNs), especially U-Net (Soibam et al., 2023; Malakhov et al., 2023), have successfully segmented two-phase flow (TPF) images by learning complex features in data, thus providing a promising solution to automate HSV PD segmentation. Unlike the shadowgraphy images typically used for TPF studies (Figure 1a), PD images feature complex bubble structures and diverse liquid-vapor interactions. These unique attributes create challenges for accurate segmentation, particularly in identifying the bubble footprints and contact lines.
@@
Figure 1: (a) Sample of Front-Lit Shadowgraphy images from two-phase flow. (b) Sample Phase-Detection Images were used in this study.
@@
HSV PD data offers a valuable lens into the boiling process by delineating liquid, vapor, and microlayer phases. However, these HSV datasets’ manual or semi-automated analysis can be labor-intensive and subject to subjectivity, limiting scalability for large multimodal datasets. Traditional segmentation techniques, including edge detection and thresholding, struggle with overlapping bubbles and dynamic flow patterns, often requiring manual adjustments across experiments. Recently, CNN architectures like U-Net have become the standard for HSV tasks due to their ability to learn intricate visual patterns and deliver precise segmentation results (Zhang et al., 2020; Eppel et al., 2020; Brzozowski et al., 2022). Nevertheless, these CNN architectures, though powerful, exhibit limited generalizability when applied to multimodal datasets comprising varying experimental conditions, posing a significant barrier to their adoption in scientific PD tasks (Seong et al., 2023; Ravichandran et al., 2023; Passoni et al., 2024).
@@
This limitation is pronounced in HSV PD analysis, where dynamic physical processes and domain-specific visual characteristics demand flexible yet robust segmentation models. The emergence of vision foundation models, like the Segment Anything Model (SAM), has demonstrated promising generalization capabilities across various segmentation tasks by leveraging transformer-based architectures (Zhang et al., 2023). Despite their potential, adapting these models to scientific HSV PD tasks remains underexplored, with most applications focused on natural images rather than PD in boiling experiments (Portillo-Portillo et al., 2022).
@@ 2 Related Works
2.1 Research Gaps
The development of image segmentation techniques for TPF analysis has shown significant progress, transitioning from classical image processing methods (Wenyin et al., 2008; Paz et al., 2017; Zhou & Niu, 2020; Richenderfer et al., 2018; Jin & Shirvan, 2021; Hanafizadeh et al., 2011; Singh et al., 2009) to advanced deep learning models (Passoni et al., 2024; Malakhov et al., 2023; Kim & Park, 2021; Soibam et al., 2023; Seong et al., 2023; Ravichandran et al., 2023; Suh et al., 2024). Despite these advancements, several critical gaps remain unaddressed, limiting the scalability and precision of current methods for PD tasks in TPF systems.
First, while CNNs such as U-Net (Ronneberger et al., 2015) have demonstrated success in segmenting simpler shadowgraphy-based TPF images (Jüngst et al., 2024; Choi et al., 2022), their potential remains entirely unexplored for the more intricate PD tasks. Similarly, scalable VFMs like SAM (Kirillov et al., 2023), despite achieving groundbreaking segmentation performance across diverse domains, have yet to be applied to either shadowgraphy-based TPF segmentation or the complex HSV PD challenges. This dual gap highlights a significant opportunity to leverage the robust feature extraction capabilities of CNNs and the scalability, adaptability, and fine-grained precision of VFMs to address real-world HSV PD segmentation complexities, unlocking transformative advancements in the field.
Moreover, there is a notable lack of open-source multimodal HSV datasets tailored specifically for PD segmentation. Existing datasets (Dunlap et al., 2022; Hu, 2023; Hassan et al., 2023) suffer from significant limitations. First, they are derived predominantly from simplistic front/back-lit shadowgraphy techniques or theoretical simulations, making them inadequate for real-world experimental PD conditions. Second, these datasets are often unimodal, limited in scale, and lacking the diversity and complexity required to develop scalable, generalizable models. Finally, this absence of representative, large-scale experimental datasets restricts the potential for developing robust VFMs tailored to PD segmentation tasks.
Furthermore, existing UQ studies (Selvan et al., 2020; Munia et al., 2025; Wahid et al., 2024; Liu et al., 2019) primarily focus on general-purpose applications, with limited attention to HSV PD analysis. Moreover, these approaches rarely incorporate real experimental metrics such as contact line density and dry area fraction at the pixel level—two key parameters for accurate and reliable TPF segmentation. This omission hinders the development of robust models capable of quantifying uncertainties inherent in PD segmentation.
Therefore, this paper builds upon prior work in (Maduabuchi et al., 2024; Maduabuchi, 2024) by highlighting the evolution of image processing techniques, the role of deep learning, and the challenges posed by existing datasets and UQ methods in Appendices A.1A.5. By leveraging scalable VFMs, robust uncertainty quantification, and newly curated open-source experimental datasets, our work establishes a transformative framework for advancing HSV PD segmentation.
2.2 Technical Contributions
To address the critical limitations in existing TPF segmentation frameworks, our work introduces a comprehensive and scalable approach tailored specifically for HSV PD segmentation. The key contributions are as follows:
2.2.1 Transfer Learning for PD Analysis
We bridge the gap in pixel-level segmentation for PD by adapting transfer learning from biological cell imaging domains to HSV datasets. This ensures that custom U-Net models can effectively capture the fine distinctions between vapor, liquid, and microlayer phases, overcoming the object-level boundary limitations of traditional CNNs like U-Net and Mask R-CNN. By leveraging domain knowledge from analogous tasks, our models achieve superior performance under complex, multimodal PD datasets encompassing diverse boiling conditions and fluctuating heat flux scenarios.
2.2.2 Integrated Uncertainty Quantification
We introduce a robust UQ pipeline that systematically quantifies errors in pixel-level segmentation metrics, including contact line density and dry area fraction. Unlike existing general-purpose UQ methods, our framework directly addresses the variability inherent in experimental PD data, ensuring reliable and interpretable outputs. Open-source scripts for quantifying discretization errors further enhance reproducibility and accuracy in assessing phase transitions and boiling heat transfer dynamics.
2.2.3 Open-Source Multimodal PD Datasets.
To overcome the scarcity of representative datasets for PD tasks, we curate and release the first large-scale, multimodal HSV datasets tailored for PD segmentation. These datasets encompass diverse experimental conditions and boiling regimes, significantly surpassing existing unimodal and simulation-driven datasets (Dunlap et al., 2022; Hu, 2023; Hassan et al., 2023). This contribution provides a foundation for developing scalable, generalizable segmentation models.
2.2.4 Extension to Vision Foundation Models
We demonstrate, for the first time, the applicability of scalable VFMs, such as SAM (Kirillov et al., 2023), to PD segmentation. By fine-tuning SAM on our curated PD datasets, we bridge the scalability gap, enabling VFMs to generalize effectively across diverse modalities and experimental conditions. This advancement ensures seamless adaptation to new PD scenarios, addressing the specificity constraints of earlier CNN-based approaches.
2.2.5 Holistic Framework for PD Analysis
Our integrated framework—spanning segmentation, uncertainty quantification, dataset development, and cross-dataset adaptability—provides an open-source, reproducible solution for advancing PD segmentation research. By equipping the community with datasets, tools, and scalable models, we establish a versatile platform for addressing the complexities of boiling heat transfer, fostering innovation across experimental and computational research contexts.